dukedadon:

Which set of line segments could create a right triangle? (4 points) Group of answer choices 15, 30, 35 15, 36, 39 15, 20, 29 5, 15, 30

Vocaloid:

Please do not respond to questions unless you know how to solve the problem. a^2 + b^2 = c^2 c will always be the hypotenuse, the largest side. for each answer choice, plug in a, b, and c and see if both sides are equal. if both sides of the equation are equal, it's a right triangle.

dukedadon:

@vocaloid wrote:
Please do not respond to questions unless you know how to solve the problem. a^2 + b^2 = c^2 c will always be the hypotenuse, the largest side. for each answer choice, plug in a, b, and c and see if both sides are equal. if both sides of the equation are equal, it's a right triangle.
thank you voca, for properly helping.